### v1

# Chiral anomaly in the Weyl semimetal TaRhTe4

Abstract
TaRhTe4 is a type-II Weyl semimetal, exhibiting four Weyl points in proximity to the Fermi level. In this article, we report our results of a systematic study of longitudinal magnetoresistance in TaRhTe4. Our findings indicate that magnetoresistance becomes negative only when the magnetic field is applied parallel to the electric field. By rotating $\mathbf{E}$ (as well as $\mathbf{B}$), we show that its origin is consistent with the prediction of the chiral anomaly, while the current jetting effect and weak localization could be excluded. The negative magnetoresistance persists up to room temperature, suggesting that TaRhTe4 exhibits distinctive properties within the family of Weyl semimetals.
Keywords: Chiral anomaly, Negative magnetoresistance, Weyl semimetals, Current jetting effect, TaRhTe4
I Introduction
In the last decade, tremendous attention has been devoted to the study of Weyl and Dirac semimetals. These are novel classes of topological materials, hosting relativistic quasiparticles named Weyl and Dirac fermions Yan and Felser (2017); Hasan et al. (2017). The electronic band structure of these materials is characterized by band-touching points, identified as Weyl and Dirac points, respectively, and linear quasiparticle dispersion in proximity to these points Armitage et al. (2018).
In 3D Dirac semimetals (DSMs) both inversion and time reversal symmetries are present, making Dirac cones four-fold degenerate. In contrast, in Weyl semimetals (WSMs), either time-reversal symmetry (TRS) or inversion symmetry (IS) is broken, which results in Weyl points (WPs) being non-degenerate and described by a topological charge. The topological nature of WPs leads to many interesting properties, including their protection against small perturbations and distortions Armitage et al. (2018). In addition, WSMs show many interesting phenomena, which can be observed experimentally. One of those is the development of protected surface states, called Fermi arcs, which can be identified through angle-resolved photoemission spectroscopy and represent the fingerprint of the Weyl cones in the bulk (Liu et al., 2016). Another effect predicted in WSMs is the chiral anomaly, which is characterized by the anomalous non-conservation of a chiral current Nielsen and Ninomiya (1983); Yang et al. (2015); Huang et al. (2015); Niemann et al. (2017a); Li et al. (2016a).
The chiral anomaly manifests itself in transport properties as a negative contribution to the longitudinal magnetoresistance (MR) when the applied magnetic field is parallel to the electric field that causes the motion of the carriers Son and Spivak (2013); Burkov (2014). Figure 1a shows a schematic representation of this effect. Such phenomenology has been observed in several Weyl semimetals, including NbP, NbAs Niemann et al. (2017b); Li et al. (2017), Dirac semimetals ZrTe5 Li et al. (2016a), Na3Bi Xiong et al. (2015), Cd3As2 Li et al. (2016b), GdPtBi Hirschberger et al. (2016); Liang et al. (2018). In all of these cases, numerous WPs are close to the Fermi level, making it difficult to attribute the effect to any specific couple of WPs.
Here, we study TaRhTe4, a sibling compound of TaIrTe4, which is the simplest example of WSMs, often referred to as the ”hydrogen atom” in the world of WSMs due to its minimal number of WPs near the Fermi level, which must be four for non-centrosymmetric WSMs Khim et al. (2016); Koepernik et al. (2016); Belopolski et al. (2017). TaIrTe4 has many interesting properties, both from the point of view of fundamental physics and from the perspective of next-generation electronics. For instance, it shows strong positive MR at low temperatures Khim et al. (2016), presents room-temperature nonlinear Hall effect Kumar et al. (2021), surface superconductivity Xing et al. (2020) and anisotropic photo-response Lai et al. (2018); Zhuo et al. (2021). TaRhTe4 is much less investigated Zhang et al. (2024).
TaRhTe4 crystallizes in the same orthorhombic non-centrosymmetric space group (No: Pmn21) structure as TaIrTe4 and it is a layered van der Waals (vdW) material. The crystal structure can be derived from WTe2 Shipunov et al. (2021), as shown in Figure 1b.
Figure 1: (a) Schematic illustration of bulk Landau levels of a pair of Weyl points. The linear lines represent the zeroth quantum Landau Level with blue and red chiralities in a $\mathbf{B}\parallel\mathbf{E}$. (b) Crystal structure of the TaRhTe4 compound.
In this article, we report our systematic investigation of the magnetotransport properties of TaRhTe4 single crystal. We measured the MR signal by applying a magnetic field up to 10 T in various geometries, including in-plane (parallel and perpendicular to the electric field) and out-of-plane (perpendicular to the electric field). Surprisingly, the magnetotransport properties of TaRhTe4 strongly differ from TaIrTe4 (shown in supplementary information Fig. S1 Behnami and et al (2024)) despite the similarity of the crystal structures and electronic band structures Shipunov et al. (2021). In particular, in the TaRhTe4 sample, we observe a negative MR when the electric and magnetic fields are parallel to each other, either oriented along the crystallographic a- and b- directions. Furthermore, this negative MR persists up to room temperature, suggesting that TaRhTe4 possesses unique properties within the family of WSMs.
II Experimental
The single crystals were grown using the self-flux method. To prepare the crystals for further study, the surfaces, which were contaminated by a small amount of flux, were mechanically cleaved Shipunov et al. (2021). The orientation of the crystal axis has been systematically determined by x-ray Laue diffraction.
The setup for DC electric transport measurements was prepared by gluing the electrodes (silver wires with 50 $\mu m$ thick) to the samples using silver paint (Dupont 4929N) as shown in the inset Figure 2a. All transport measurements were carried out within a temperature range of 8 K to room temperature, using a tested homemade probe and Oxford cryostats equipped with sufficiently large magnets. A Keithley Standard Series 2400 Source Measure Unit (SMU) was used to apply the electric current (field), while a Keithley Nanovoltmeter Model 2182A was utilized to measure the MR signal.
In each experiment, we measured the longitudinal resistivity signal. To investigate the anisotropy of the samples, we studied the total magnetic field dependence up to 10 T for two magnetic field geometries: in-plane (parallel and perpendicular to the electric field) and out-of-plane (perpendicular to the electric field).
III Results
Figure 2 shows the magnetotransport measurements for the first sample of TaRhTe4. More in detail, Figure 2a presents the temperature dependence of the electrical resistivity in a zero magnetic field. The curve shows a metallic behavior decreasing monotonically with the temperature and a moderate RRR= 2.23. Figures 2b, c, and d show the dependence of resistivity on the magnetic field. In this experiment, the electric current was injected along the crystallographic $a$-axis and, as a consequence, the electric field results oriented along the same axis. The magnetic field was applied sequentially along all three crystallographic axis, as shown in the inset of Figure 2a. The presented MR is defined as MR $=100\times\frac{\rho\,(\mathbf{B})-\rho\,(0)}{\rho\,(0)}$, where $\rho\,(\mathbf{\mathbf{B}})$ and $\rho\,(0)$ are the resistivities of the sample in a magnetic field $\mathbf{B}$ and in zero magnetic field, respectively.
[…]
Figures 2c and 2d present the magnetoresistance as a function of the magnetic field applied along the crystal $b$- and $c$-axis, respectively. In both cases, MR is always positive and relatively small, reaching the maximum value of a few percent. In addition, it presents a parabolic behavior, which is indicative of the cyclotronic origin of the magnetoresistive effect Li et al. (2016b).
Figure 2: Longitudinal magnetoresistance (MR) in the first sample of TaRhTe4. (a) Temperature dependence of resistivity. Inset: a microscopic image of our first TaRhTe4 single crystal was prepared for electric measurement. (b) Longitudinal MR in the case of an in-plane magnetic field with $\mathbf{B}\parallel\mathbf{E}$. (c) Longitudinal MR in the case of an in-plane magnetic field with $\mathbf{B}\perp\mathbf{E}$. (d) Longitudinal MR in the case of out-of-plane magnetic field with $\mathbf{B}\perp\mathbf{E}$.
In order to confirm the negative MR, observed in Figure  2b, we deeply investigated a second sample. To this aim, we chose a slab-shaped sample of TaRhTe4 and we attached six electrodes on top of that, which enables us to measure the MR signal along different crystallography axis (see inset of Figure 3). We first measured the zero-field resistivity of the second sample in different configurations, as shown in Figure 3. Interestingly, the resistivity estimated with voltage electrodes along the $a$-axis (V26 and V35) and current prevalently injected along the $a$-axis is relatively different with respect to the resistivity evaluated putting both voltage and current electrodes along the sample $b$-axis (I34 and V56). The different RRR values (2.68, 2.55, and 1.44) indicate the relative anisotropy between the $a$-axis and the $b$-axis conductivities. From the finite element simulations (see supplementary information Fig. S2) and from the comparison of the experimental values of zero-field resistances measured in the three configurations, we extracted the value of in-plane resistivity anisotropy $\rho_{b}/\rho_{a}$=6.8.
After a zero-field characterization, we measured the sample magnetoresistance. For this specimen, we explored the following configurations.
Figure 3: Temperature dependence of resistivity for three configurations of the second sample of TaRhTe4. Inset: A microscopic picture of the second sample of TaRhTe4. The numbers indicate the electrodes that we used to apply the electrical field and measure longitudinal resistivity.
Magnetic field in the ab-plane and electric field parallel to the $a$-axis of the crystal.
In the first setup, we injected current from contact (1) to contact (5) and measured the voltage signal between contact (2) and contact (6), with the magnetic field applied parallel to the $a$-axis of the crystal. As illustrated in Figure 4a, we observed a negative MR, which reaches the maximum value of 0.22 % and persists up to high temperatures. In the second setup, we changed the direction of the magnetic field from parallel to perpendicular to the $a$-axis. This time, we observed a positive MR which reaches 3.9 % for $T=9$ K and $\mathbf{B}=10$ T, as depicted in Figure 4d.
Magnetic field in the ab-plane and electric field parallel to the $b$-axis of the crystal.
In this configuration, we injected current from contact (3) to contact (4) and measured the voltage signal between contact (5) and contact (6). Firstly, we applied a magnetic field parallel to the $b$-axis of the crystal, observing again a negative MR (up to 0.19 %) as illustrated in Figure 4b. Secondly, we changed the direction of the magnetic field from parallel to perpendicular to the sample $b$-axis. This time, we observed a positive MR up to 4.2 % at $T=9$ K and $\mathbf{B}=10$ T, as depicted in Figure 4e.
Magnetic field in the ab-plane and electric field at 30 degrees from $a$-axis of the crystal:
In the third configuration, we injected current from contact (1) to contact (6) and measured the voltage signal between contact (3) and contact (5). In the first experiment, we applied a magnetic field parallel to both an electric field and the diagonal axis of the sample, tilted about 30 degrees from the $a$-axis. Clearly, such an angle is only a nominal value, not necessarily corresponding to the exact direction of the electric field, which strictly depends on the actual current path. As illustrated in Figure 4c, we observed a positive MR. At $T=8$ K and $\mathbf{B}=10$ T, the highest positive MR of 0.45 % can be obtained. In the second configuration, we changed the direction of the magnetic field from parallel to perpendicular and measured the signal. Once again, we observed a positive MR, as depicted in Figure 4f. At $T=9$ K and $\mathbf{B}=10$ T, a positive MR of 3.2 % is measured. Therefore, we did not observe negative MR in this setup, whether the magnetic field was parallel or perpendicular.
Figure 4: Longitudinal MR in the second sample of TaRhTe4. (a-c) Longitudinal MR in the case of an in-plane magnetic field with $\mathbf{B}\parallel\mathbf{E}$. The applied magnetic field is along $a$-, $b$-, and 30 degrees from the $a$-axis of the crystal respectively. (d-f) Longitudinal MR in the case of an in-plane magnetic field with $\mathbf{B}\perp\mathbf{E}$. The applied magnetic field is along $b$-, $a$-, and 30 degrees from the $b$-axis of the crystal respectively.
IV Discussion
In non-magnetic conductors, the presence of external magnetic fields typically generates a magnetoresistive effect caused by the cyclotronic motion of the charge carriers, when the field itself has a component perpendicular to the direction of the current density vector Li et al. (2016b). Such semiclassical effect strongly depends on both intrinsic and extrinsic properties of the measured sample, such as the effective mass of the charge carriers and the scattering time, and, in the low field limit, is proportional to B2. In general, clean systems exhibit higher cyclotronic magnetoresistance, intuitively depending on the number of cyclotronic orbits traveled by charge carriers between a scattering event and the following one. As expected, in our case, this is also the dominant contribution in all the configurations with the magnetic and electric fields not aligned in the same direction.
In Weyl semimetals, an additional negative contribution to the magnetoresistance is possibly caused by the chiral anomaly. Equivalently, the chiral anomaly manifests itself as a positive contribution to the electrical conductivity in magnetic field Son and Spivak (2013); Burkov (2016), only when it is parallel to the electric field that causes the motion of charge carriers, namely to the direction of the injected current. In this situation, the correction to the electrical conductivity due to the chiral anomaly for a type-I Weyl semimetal reads Son and Spivak (2013):
$\Delta\sigma(B)=\frac{e^{2}}{4\pi^{2}\hbar c}\frac{v_{F}}{c}\frac{(eBv_{F})^{2}}{\epsilon_{F}^{2}}\tau_{a},$ (1)
where $\tau^{-1}_{a}$ is the longitudinal axial current relaxation rate, $c$ is the light speed and $\epsilon_{F}$ and $v_{F}$ are the Fermi energy and the Fermi velocity respectively. Remarkably, it has been demonstrated that the quadratic-in-field dependence holds also in case of a type-II, or tilted WSMs, for which a positive longitudinal magneto-conductivity appears for all finite angles between the tilt direction of the cones and the direction of the applied magnetic/electric field Sharma et al. (2017). Figure 5 presents the positive magnetoconductivity of the first sample of TaRhTe4 in the experimental configuration of Figure 2b. All the curves are nicely reproduced by the $\mathbf{B}^{2}$ fitting function, corroborating the chiral anomaly as a possible source for the observed effect. Here, $\Delta\sigma=\sigma(\mathbf{B})-\sigma(\mathbf{B}=0)$, where $\sigma(\mathbf{B})$ and $\sigma(\mathbf{B=0})$ are the conductivity of the first sample in a magnetic field and in zero magnetic fields respectively. The same result is obtained for the second sample as shown in supplementary information Figures S3 and S4 Behnami and et al (2024).
Figure 5: Positive contribution of electrical conductivity as a function of magnetic field for the first sample of TaRhTe4. The direction of a magnetic field is parallel to the electric field ($\mathbf{B}\parallel\mathbf{E}$).
In the following, we will briefly discuss other possible sources of negative magnetoresistance in non-magnetic materials. The first possible phenomenon that can cause negative MR is the weak localization effect Altshuler et al. (1980); Lee and Ramakrishnan (1985); Qu et al. (2024).However, weak localization is typically observed in 2D or highly anisotropic (out-of-plane/in-plane anisotropy of electronic properties) materials, when the magnetic field is applied in the out-of-plane direction. In this condition, the threading of magnetic flux in interfering current paths may result in negative magnetoresistance. In our case, the negative MR appears when field and current are both in-plane and parallel to each other, which is the configuration that minimizes the threading of magnetic flux in interfering current paths. Moreover, both the magnitude and the magnetic field dependence of our magnetoresistance are not described by conventional models for weak localization (see Supplemental InformationBehnami and et al (2024)).
The main alternative to the chiral anomaly to explain the observed phenomenology is represented by the so-called current jetting effect, which appears as a semiclassical negative longitudinal MR caused by a non-uniform current distribution in the sample when $\mathbf{B}\parallel\mathbf{E}$. Such an effect typically appears when current contacts are smaller than the cross-section of the sample, causing field-induced steering of the current to the direction of the magnetic field Liang et al. (2018); Sykora et al. (2020). In the measurement configuration of Figure 2b, referring to sample one, the size of the current electrodes is comparable with the sample width (see inset of Figure 2a). Therefore, we do not expect any substantial current jetting. Remarkably, in the case of sample two, where the electrodes are smaller and differently distributed across the sample (see inset of Figure 3), the experimental configurations in Figure 4a and 4b returns almost the same value of negative magnetoresistance, very similar to the one observed in sample one. In addition, we tested the influence of size and position of the electrical contacts on the appearance of the negative MR by measuring several other samples. In all the experiments we verified that the only relevant parameter is the mutual orientation of current, magnetic field and crystal axis, which have to be parallel to each other, as expected in case of chiral anomaly (see Supplementary InformationBehnami and et al (2024)).
Finally, we note that, in sample one, at $T<$100 K the negative magnetoresistive curves of Figure 2b change their concavity, with a tendency to turn up at high fields. This effect can be trivially explained through an involuntary tilting of the applied magnetic field towards the out-of-plane direction, which causes an additional positive cyclotronic component that becomes dominant for sufficiently large fields. This spurious effect also accounts for the non monotonic trend in temperature of the curves in Figure 2b. The curves a $T$=200 and 250 K do not show the upturn due to the strong suppression of the cyclotronic magnetoresitance at high temperatures. Hence, the insensitivity of the observed magnetoresistance to the geometry and distribution of the current electrodes on the sample excludes with reasonable certainty the current jetting as a possible source for the negative magnetoresistance, in favor of an intrinsic phenomenon related to its non-trivial topology.
V Conclusions
In conclusion, we have systematically measured the MR of two single crystals of TaRhTe4, with different experimental configurations by varying both the distribution of the current/voltage electrodes across the samples and the applied magnetic field direction. In the setups where the applied electric and magnetic fields are parallel to each other and aligned either along the $a$- or the $b$- axis, we always measured a negative magnetoresistance (or positive magnetoconductance) compatible with a chiral anomaly effect. The robustness of the observed phenomenology with respect to the change of geometry in the current electrodes allowed us to exclude the current jetting as an alternative source for the effect.
Supporting Information
Supporting Information is available from the Online Library or from the author Behnami and et al (2024).
Acknowledgements
We would like to thank Tino Schreiner and Danny Baumann for their technical support. MB acknowledges Forschungsstipendien für Doktorandinnen und Doktoranden, 2024 (DAAD) via project number 57694190 and the IFW Excellence Program 2018. SA acknowledges DFG via project number 523/4-1. We would like to acknowledge the financial support provided by the Deutsche Forschungsgemeinschaft (DFG) through Project number 449494427. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Wurzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter—ct.qmat (EXC 2147, project ID 242021). The project was supported by the Leibniz Association through the Leibniz Competition.
Conflict of Interest
The authors declare no conflict of interest.
Author Contributions
B.B., D.V.E., H.R., F.C., and M.B. proposed the study. G.Sh., B.R.P., C.G.F.B. and S.A. prepared and characterized the samples. M.B., H.R., V.K., M.P., J.D., I.P., D.V.E., and F.C. designed the measurement setup, performed the experiments, and analyzed the data. I.P. performed the finite element simulations. B.B., H.R., and F.C. supervised the study. M.B., H.R., D.V.E., and F.C. wrote the manuscript with input from all authors.